### v1

# Strong Converse Exponent of Quantum Dichotomies

Abstract
The quantum dichotomies problem asks at what rate one pair of quantum states can be approximately mapped into another pair of quantum states. In the many copy limit and for vanishing error, the optimal rate is known to be given by the ratio of the respective quantum relative distances. Here, we study the large-deviation behavior of quantum dichotomies and determine the exact strong converse exponent based on the purified distance. Our result is characterized by a simple optimization of quantum Rényi information measures involving all four mutually non-commuting quantum states. Our findings thus constitute an operational example of studying multivariate extensions of quantum Rényi information measure.
I Introduction
Quantum Rényi information measures are basic tools in quantum information theory. Previous measures are usually defined for two non-commuting quantum states, but currently there is an increasing interest in finding quantum Rényi information measure with more than two arguments. The recent works MBV2024geometric ; VTH2024matrix ; FFHT2024matrix have made some progress in this topic (see also BunthVrana2022equivariant ; FLO2023monotonic ; MNW2024on ; NMLW2024optimal ).
One method to find meaningful quantum Rényi information measures is to study the large deviation behaviors of operational quantum information processing tasks. The quantum Rényi information measures usually appear as the large deviation-type exponential rate. In this paper, we study quantum dichotomies transformations. Consider two quantum dichotomies $(\rho_{1},\sigma_{1})$ and $(\rho_{2},\sigma_{2})$, quantum dichotomies transformation is a procedure of using a quantum channel to transform $(\rho_{1},\sigma_{1})$ into as much $(\rho_{2},\sigma_{2})$ as possible, in which we require that the transformation $\sigma_{1}\rightarrow\sigma_{2}$ is exact and the transformation $\rho_{1}\rightarrow\rho_{2}$ is allowed to be approximate FDT2019information ; WangWilde2019resource . It is an important quantum information processing task which have wide variety of applications in quantum resource theory FDT2019information ; WangWilde2019resource ; LPCRTK2024quantum ; KCT2019avoid ; KCT2018Beyond ; KCT2019Moderate ; KumagaiHayashi2013second ; Wilde2022distinguishability . Because there exists four mutually non-commuting quantum states in this problem, it is a proper task for finding multi-variate quantum Rényi information measures. The commonly used distances to measure the closeness between the final state and the target state $\rho_{2}$ are the trace distance and the purified distance. In the asymptotic setting when unlimited $(\rho_{1},\sigma_{1})$ are available, no matter which distance we choose, the optimal rate at which the asymptotically perfect transformation can be achieved equals to the ratio between the quantum relative entropy $D(\rho_{1}\|\sigma_{1})$ and $D(\rho_{2}\|\sigma_{2})$ FDT2019information ; WangWilde2019resource .
Although the first-order asymptotic analysis of quantum dichotomies transformation has been well understood, the finer asymptotic results, including the second-order expansion, moderate-deviation analysis and large-deviation type exponential analysis, are more complicated and they depend on the choice of the distance. The reference LPCRTK2024quantum established the exact second-order asymptotic, moderate-deviation expansion and large-deviation type exponential behavior based on the trace distance under the assumption $\rho_{2}$ commutes with $\sigma_{2}$. The reference KCT2018Beyond and KCT2019Moderate established the second-order asymptotic and the moderate-deviation analysis based on the purified distance, respectively, when $(\rho_{1},\sigma_{1})$ and $(\rho_{2},\sigma_{2})$ are both classical dichotomies. To the best of our knowledge, the current research on the finer asymptotic analysis of quantum dichotomies transformation are only restricted in classical or semi-classical setting. The related results are still unknown in fully quantum setting.
In this paper, we study the high-error large-deviation exponential behavior for quantum dichotomies transformation, also known as the strong converse exponent for this task, which is the best rate of exponential convergence of the performance of this task towards the worst. We prove that for any quantum dichotomies $(\rho_{1},\sigma_{1})$ and $(\rho_{2},\sigma_{2})$, when the transformation rate $r$ is larger than the first-order asymptotic $D(\rho_{1}\|\sigma_{1})\cdot D(\rho_{2}\|\sigma_{2})^{-1}$, the strong converse exponent based on the purified distance is given by
$\sup_{\frac{1}{2}\leq\alpha\leq 1}\frac{1-\alpha}{\alpha}\Big{\{}rD_{\alpha}^{*}(\rho_{2}\|\sigma_{2})-D_{\alpha/(2\alpha-1)}^{*}(\rho_{1}\|\sigma_{1})\Big{\}},$ (1)
where $D_{\alpha}^{*}(\rho\|\sigma)$ is the sandwiched Rényi divergence. Our work is the first time to establish the exact strong converse exponent for this task in fully quantum setting, filling the gap in previous work. Our strong converse exponent is characterized by a quantum Rényi information measure involving four quantum states.
In addition, if $\rho_{2}$ is a pure state, we also show that the strong converse exponent based on the trace distance is
$\sup_{\beta\geq 1}\frac{\beta-1}{\beta}\Big{\{}rD_{1/\beta}(\rho_{2}\|\sigma_{2})-D_{\beta}^{*}(\rho_{1}\|\sigma_{1})\Big{\}},$ (2)
where $D_{\alpha}(\rho\|\sigma)$ is the Petz Rényi divergence. We conjecture that this formula still holds for general quantum states $\rho_{2}$. The Petz Rényi divergence was used to characterize the reliability functions (the best rate of exponential convergence of the error of a task towards the perfect) for quantum information processing tasks in previous research AKCMBMA2007discriminating ; NussbaumSzkola2009chernoff ; Li2016discriminating ; Nagaoka2006converse ; Hayashi2007error ; ANSV2008asymptotic ; Renes2022achievable ; CHDH2020non ; HayashiTomamichel2016correlation . If the conjecture is correct, this will provide new type of operational meaning for the Petz Rényi divergence.
In the derivation of the achievability part of the strong converse exponent based on the purified distance, we use a technique developed by Mosonyi and Ogawa MosonyiOgawa2017strong , to derive a weaker bound characterized by the log-Euclidean Rényi divergence and then improve it to the final bound, one of the authors also use the similar technique to establish the strong converse exponents for privacy amplification and quantum information decoupling LiYao2024operational . The proof of the optimality part comes from an application of the operator Hölder inequality essentially. For the strong converse exponent based on the trace distance, we mainly make use of an improved Fuchs-van de Graaf to relate it to the strong converse exponent based on the purified distance.
The remainder of this paper is organized as follows. In Section II, we introduce some preliminaries in quantum information theory. In Section III, we present the main problem and the main result. In Section IV, we give the proof of the achievability part of the strong converse exponent for quantum dichotomies transformation. In Section V, the corresponding proof of the optimality part is presented. In Section VI, we conclude this paper with some discussion.
II Preliminaries
II.1 Notation
Let $\mathcal{H}$ be a finite-dimensional Hilbert space, we denote the set of linear operators on $\mathcal{H}$ as $\mathcal{L}(\mathcal{H})$. Let $\mathcal{P}(\mathcal{H})$ and $\mathcal{S}(\mathcal{H})$ represent the set of positive semi-definite operators and the set of quantum states, respectively. The support of an operator $A\in\mathcal{L}(\mathcal{H})$ is denoted as ${\operatorname{supp}}(A)$. For $\rho\in\mathcal{S}(\mathcal{H})$, we use the notation $\mathcal{S}_{\rho}$ for the set of quantum states whose supports are contained in ${\operatorname{supp}}(\rho)$. We denote as $|\mathcal{H}|$ the dimension of $\mathcal{H}$.
The trace distance and purified distance are two most commonly used distances to measure the closeness of two quantum states. For $\rho,\sigma\in\mathcal{S}(\mathcal{H})$, the trace distance $d(\rho,\sigma)$ and the purified distance $P(\rho,\sigma)$ are given by
$\begin{split}&d(\rho,\sigma):=\frac{\|\rho-\sigma\|_{1}}{2},\\ &P(\rho,\sigma):=\sqrt{1-F^{2}(\rho,\sigma)},\end{split}$ (3)
where $\|\cdot\|_{1}$ is the Schatten-1 norm and $F(\rho,\sigma):=\|\sqrt{\rho}\sqrt{\sigma}\|_{1}$ is the fidelity function.
A quantum channel $\mathcal{N}$ is a completely positive and trace-preserving map. Let $A$ be a self-adjoint operator with spectral projections $P_{1},\ldots,P_{r}$. The pinching channel $\mathcal{E}_{A}$ associated with $A$ is defined as
$\mathcal{E}_{A}:X\mapsto\sum_{i=1}^{r}P_{i}XP_{i}.$
The pinching inequality Hayashi2002optimal tells that for any $\sigma\in\mathcal{P}(\mathcal{H})$, we have
$\sigma\leq v(A)\mathcal{E}_{A}(\sigma),$ (4)
where $v(A)$ is the number of different eigenvalues of $A$.
II.2 Quantum Rényi divergences
The classical Rényi divergence Renyi1961measures is an information quantity which play important roles in many fields. Because of the non-commutativity of quantum states, there are infinite in-equivalent quantum generalizations of classical Rényi divergence. In this section, we introduce the sandwiched Rényi divergence MDSFT2013on ; WWY2014strong , the Petz Rényi divergence Petz1986quasi and the log-Euclidean Rényi divergence HiaiPetz1993the ; MosonyiOgawa2017strong which have been found wide applications so far AKCMBMA2007discriminating ; NussbaumSzkola2009chernoff ; Li2016discriminating ; Nagaoka2006converse ; Hayashi2007error ; ANSV2008asymptotic ; Renes2022achievable ; CHDH2020non ; HayashiTomamichel2016correlation ; MosonyiOgawa2015quantum ; GuptaWilde2015multiplicativity ; CMW2016strong ; LiYao2024strong ; LYH2023tight ; LiYao2024reliability ; Hayashi2015precise ; LiYao2021reliable .
Definition 1
Let $\alpha\in(0,+\infty)\setminus\{1\}$, $\rho\in\mathcal{S}(\mathcal{H})$ and $\sigma\in\mathcal{P}(\mathcal{H})$. When $\alpha>1$ and ${\operatorname{supp}}(\rho)\subseteq{\operatorname{supp}}(\sigma)$ or $\alpha\in(0,1)$ and ${\operatorname{supp}}(\rho)\not\perp{\operatorname{supp}}(\sigma)$, the sandwiched Rényi divergence and the Petz Rényi divergence of order $\alpha$ are respectively defined as
$\begin{split}&D_{\alpha}^{*}(\rho\|\sigma):=\frac{1}{\alpha-1}\log Q_{\alpha}^{*}(\rho\|\sigma),\quad\text{with}\ \ Q_{\alpha}^{*}(\rho\|\sigma)=\operatorname{Tr}{({\sigma}^{\frac{1-\alpha}{2\alpha}}\rho{\sigma}^{\frac{1-\alpha}{2\alpha}})}^{\alpha},\\ &D_{\alpha}(\rho\|\sigma):=\frac{1}{\alpha-1}\log Q_{\alpha}(\rho\|\sigma),\quad\text{with}\ \ Q_{\alpha}(\rho\|\sigma)=\operatorname{Tr}\rho^{\alpha}\sigma^{1-\alpha};\end{split}$ (5)
otherwise, we set $D_{\alpha}^{*}(\rho\|\sigma)=D(\rho\|\sigma)=+\infty$. When $\alpha>1$ and ${\operatorname{supp}}(\rho)\subseteq{\operatorname{supp}}(\sigma)$ or $\alpha\in(0,1)$ and ${\operatorname{supp}}(\rho)\cap{\operatorname{supp}}(\sigma)\neq\{0\}$, the log-Euclidean Rényi divergence of order $\alpha$ is defined as
$D_{\alpha}^{\flat}(\rho\|\sigma):=\frac{1}{\alpha-1}\log Q_{\alpha}^{\flat}(\rho\|\sigma),\quad\text{with}\ \ Q_{\alpha}^{\flat}(\rho\|\sigma)=\operatorname{Tr}2^{\alpha\log\rho+(1-\alpha)\log\sigma};$ (6)
otherwise, we set $D_{\alpha}^{\flat}(\rho\|\sigma)=+\infty$.
When $\alpha\rightarrow 1$, all above quantum Rényi divergences converge to the Umegaki relative entropy Umegaki1954conditional
$D(\rho\|\sigma):=\begin{cases}\operatorname{Tr}\rho(\log\rho-\log\sigma)&\text{ if }{\operatorname{supp}}(\rho)\subseteq{\operatorname{supp}}(\sigma),\\ +\infty&\text{ otherwise.}\end{cases}$ (7)
In the following proposition, we collect some important properties of these quantum Rényi divergences.
Proposition 2
Let $\rho\in\mathcal{S}(\mathcal{H})$ and $\sigma\in\mathcal{P}(\mathcal{H})$. The sandwiched Rényi divergence, the Petz Rényi divergence and the log-Euclidean Rényi divergence satisfy the following properties.
  1. (i)
    Monotonicity in Rényi parameter MDSFT2013on ; Beigi2013sandwiched ; MosonyiOgawa2017strong : if $0\leq\alpha\leq\beta$, then $D_{\alpha}^{(t)}(\rho\|\sigma)\leq D^{(t)}_{\beta}(\rho\|\sigma)$, for $(t)=*$, $(t)=\{\}$ and $(t)=\flat$;
  2. (ii)
    Monotonicity in $\sigma$ MDSFT2013on ; MosonyiOgawa2017strong : if $\sigma^{\prime}\geq\sigma$, then $D_{\alpha}^{(t)}(\rho\|\sigma^{\prime})\leq D_{\alpha}^{(t)}(\rho\|\sigma)$, for $(t)=*$, $\alpha\in[\frac{1}{2},+\infty)$, for $(t)=\{\}$, $\alpha\in[0,2]$ and for $(t)=\flat$, $\alpha\in[0,+\infty)$;
  3. (iii)
    Variational representation MosonyiOgawa2017strong : the log-Euclidean Rényi divergence has the following variational representation
    $D_{\alpha}^{\flat}(\rho\|\sigma)=\begin{cases}\min\limits_{\tau\in\mathcal{S}(\mathcal{H})}\big{\{}D(\tau\|\sigma)-\frac{\alpha}{\alpha-1}D(\tau\|\rho)\big{\}},&\alpha\in(0,1),\\ \max\limits_{\tau\in\mathcal{S}_{\rho}(\mathcal{H})}\big{\{}D(\tau\|\sigma)-\frac{\alpha}{\alpha-1}D(\tau\|\rho)\big{\}},&\alpha\in(1,+\infty);\end{cases}$ (8)
  4. (iv)
    Data processing inequality FrankLieb2013monotonicity ; Beigi2013sandwiched ; MDSFT2013on ; WWY2014strong ; MosonyiOgawa2017strong : letting $\mathcal{N}$ be a CPTP map from $\mathcal{L}(\mathcal{H})$ to $\mathcal{L}(\mathcal{H}^{\prime})$, we have
    $D_{\alpha}^{(\rm{t})}(\mathcal{N}(\rho)\|\mathcal{N}(\sigma))\leq D_{\alpha}^{(\rm{t})}(\rho\|\sigma),$ (9)
    for $(t)=\flat$, $\alpha\in[0,1]$, for $(t)=\{\}$, $\alpha\in[0,2]$ and for $(t)=*$, $\alpha\in[\frac{1}{2},+\infty)$;
  5. (v)
    Approximation by pinching HayashiTomamichel2016correlation ; MosonyiOgawa2015quantum : for any $\alpha\geq 0$, we have
    $D_{\alpha}^{*}(\mathcal{E}_{\sigma}(\rho)\|\sigma)\leq D_{\alpha}^{*}(\rho\|\sigma)\leq D_{\alpha}^{*}(\mathcal{E}_{\sigma}(\rho)\|\sigma)+2\log v(\sigma).$ (10)
III Problem formulation and main result
Let $(\rho_{1},\sigma_{1})$ and $(\rho_{2},\sigma_{2})$ be two quantum dichotomies. In the procedure of quantum dichotomies transformation, we use a quantum channel $\mathcal{E}$ to transform $(\rho_{1},\sigma_{1})$ into as much $(\rho_{2},\sigma_{2})$ as possible. We allow a small error in the transformation $\rho_{1}\rightarrow\rho_{2}$, but the transformation $\sigma_{1}\rightarrow\sigma_{2}$ is required to be exact. For $n\in\mathbb{N}$ and $\epsilon\in[0,1]$, we define the maximal achievable quantum dichotomies transformation number within error $\epsilon$ for $n$ copies of $(\rho_{1},\sigma_{1})$ as
$M^{\Delta,\epsilon}_{\rho_{1},\sigma_{1}\rightarrow\rho_{2},\sigma_{2}}(n)$
$\quad:=\max\{m~{}:~{}\exists~{}\text{a quantum channel}~{}\mathcal{E}~{}\text{such that}~{}\mathcal{E}(\sigma^{\otimes n}_{1})=\sigma^{\otimes m}_{2},\Delta(\mathcal{E}(\rho^{\otimes n}_{1}),\rho^{\otimes m}_{2})\leq\epsilon\},$ (11)
where $\Delta$ can be chosen as the trace distance or the purified distance.
In the asymptotic setting in which multiple copies of $(\rho_{1},\sigma_{1})$ are available, the first order asymptotic behavior of $M^{\Delta,\epsilon}_{\rho_{1},\sigma_{1}\rightarrow\rho_{2},\sigma_{2}}(n)$ has been derived in FDT2019information ; WangWilde2019resource , i.e.,
$\lim_{n\rightarrow\infty}\frac{1}{n}M^{\Delta,\epsilon}_{\rho_{1},\sigma_{1}\rightarrow\rho_{2},\sigma_{2}}(n)=\frac{D(\rho_{1}\|\sigma_{1})}{D(\rho_{2}\|\sigma_{2})},$ (12)
where $\Delta$ can be either trace distance or purified distance.
Eq. (12) implies that when transformation rate $r>\frac{D(\rho_{1}\|\sigma_{1})}{D(\rho_{2}\|\sigma_{2})}$, the optimal error among all transformation schemes
$\epsilon^{\Delta,r}_{\rho_{1},\sigma_{1}\rightarrow\rho_{2},\sigma_{2}}(n):=\min\{\Delta(\mathcal{E}_{n}(\rho^{\otimes n}_{1}),\rho_{2}^{\otimes nr})~{}:~{}\exists~{}\text{a quantum channel}~{}\mathcal{E}_{n}~{}\text{such that}~{}\mathcal{E}_{n}(\sigma^{\otimes n}_{1})=\sigma^{\otimes nr}_{2}\}$
converge to $1$ exponentially fast. The exact exponential decay rate is called as the strong converse exponent of quantum dichotomies and formally defined as follows.
Definition 3
For two quantum dichotomies $(\rho_{1},\sigma_{1})$, $(\rho_{2},\sigma_{2})$ and a transformation rate $r$, the strong converse exponents of quantum dichotomies based on the purified distance and the trace distance are defined respectively as
$\begin{split}E^{P}_{sc}(\rho_{1},\sigma_{1},\rho_{2},\sigma_{2},r):&=\lim_{n\rightarrow\infty}\frac{-1}{n}\log\big{(}1-\epsilon^{P,r}_{\rho_{1},\sigma_{1}\rightarrow\rho_{2},\sigma_{2}}(n)\big{)},\\ E^{d}_{sc}(\rho_{1},\sigma_{1},\rho_{2},\sigma_{2},r):&=\lim_{n\rightarrow\infty}\frac{-1}{n}\log\big{(}1-\epsilon^{d,r}_{\rho_{1},\sigma_{1}\rightarrow\rho_{2},\sigma_{2}}(n)\big{)}.\end{split}$
Remark 4
By the relation between the purified distance and the fidelity function, $E^{P}_{sc}(\rho_{1},\sigma_{1},\rho_{2},\sigma_{2},r)$ can be equivalently expressed as
$E^{P}_{sc}(\rho_{1},\sigma_{1},\rho_{2},\sigma_{2},r)=\lim_{n\rightarrow\infty}\frac{-1}{n}\log\max_{\mathcal{E}_{n}:\mathcal{E}_{n}(\sigma_{1}^{\otimes n})=\sigma_{2}^{\otimes nr}}F^{2}(\mathcal{E}_{n}(\rho_{1}^{\otimes n}),\rho_{2}^{\otimes nr}).$ (13)
In this paper, we derive the exact expression for $E^{P}_{sc}(\rho_{1},\sigma_{1},\rho_{2},\sigma_{2},r)$ and if $\rho_{2}$ is a pure state, we also determine the exact expression for $E^{d}_{sc}(\rho_{1},\sigma_{1},\rho_{2},\sigma_{2},r)$. Our main results are as follows.
Theorem 5
For any $\rho_{1},\sigma_{1}\in\mathcal{S}(\mathcal{H}_{1})$, $\rho_{2},\sigma_{2}\in\mathcal{S}(\mathcal{H}_{2})$ satisfy ${\operatorname{supp}}(\rho_{1})\subseteq{\operatorname{supp}}(\sigma_{1})$ or ${\operatorname{supp}}(\rho_{2})\subseteq{\operatorname{supp}}(\sigma_{2})$ and any $r>0$, we have
$E^{P}_{sc}(\rho_{1},\sigma_{1},\rho_{2},\sigma_{2},r)=\sup_{\frac{1}{2}\leq\alpha\leq 1}\frac{1-\alpha}{\alpha}\big{\{}rD_{\alpha}^{*}(\rho_{2}\|\sigma_{2})-D_{\frac{\alpha}{2\alpha-1}}^{*}(\rho_{1}\|\sigma_{1})\big{\}}.$ (14)
In addition, if $\rho_{2}$ is a pure state, we have
$E^{d}_{sc}(\rho_{1},\sigma_{1},\rho_{2},\sigma_{2},r)=\sup_{\beta\geq 1}\frac{\beta-1}{\beta}\big{\{}rD_{\frac{1}{\beta}}(\rho_{2}\|\sigma_{2})-D^{*}_{\beta}(\rho_{1}\|\sigma_{1})\big{\}}.$ (15)
Remark 6
By Lemma 12 in Appendix, if $\rho_{2}$ is a pure state, we have
$\epsilon^{d,r}_{\rho_{1},\sigma_{1}\rightarrow\rho_{2},\sigma_{2}}(n)\leq\epsilon^{P,r}_{\rho_{1},\sigma_{1}\rightarrow\rho_{2},\sigma_{2}}(n)\leq\sqrt{\epsilon^{d,r}_{\rho_{1},\sigma_{1}\rightarrow\rho_{2},\sigma_{2}}(n)}.$ (16)
Eq. (16) implies that
$1-\epsilon^{d,r}_{\rho_{1},\sigma_{1}\rightarrow\rho_{2},\sigma_{2}}(n)\geq 1-\epsilon^{P,r}_{\rho_{1},\sigma_{1}\rightarrow\rho_{2},\sigma_{2}}(n)\geq\frac{1-\epsilon^{d,r}_{\rho_{1},\sigma_{1}\rightarrow\rho_{2},\sigma_{2}}(n)}{1+\sqrt{\epsilon^{d,r}_{\rho_{1},\sigma_{1}\rightarrow\rho_{2},\sigma_{2}}(n)}}.$ (17)
According to the definition of $E^{d}_{sc}(\rho_{1},\sigma_{1},\rho_{2},\sigma_{2},r)$ and $E^{P}_{sc}(\rho_{1},\sigma_{1},\rho_{2},\sigma_{2},r)$, we have
$\begin{split}&E^{d}_{sc}(\rho_{1},\sigma_{1},\rho_{2},\sigma_{2},r)\\ =&E^{P}_{sc}(\rho_{1},\sigma_{1},\rho_{2},\sigma_{2},r)\\ =&\sup_{\frac{1}{2}\leq\alpha\leq 1}\frac{1-\alpha}{\alpha}\big{\{}rD_{\alpha}^{*}(\rho_{2}\|\sigma_{2})-D_{\frac{\alpha}{2\alpha-1}}^{*}(\rho_{1}\|\sigma_{1})\big{\}}\\ =&\sup_{\frac{1}{2}\leq\alpha\leq 1}\frac{1-\alpha}{\alpha}\big{\{}rD_{2-\frac{1}{\alpha}}(\rho_{2}\|\sigma_{2})-D_{\frac{\alpha}{2\alpha-1}}^{*}(\rho_{1}\|\sigma_{1})\big{\}}\\ =&\sup_{\beta\geq 1}\frac{\beta-1}{\beta}\big{\{}rD_{\frac{1}{\beta}}(\rho_{2}\|\sigma_{2})-D_{\beta}^{*}(\rho_{1}\|\sigma_{1})\big{\}},\end{split}$ (18)
where the fourth line is because that when $\rho_{2}$ is a pure state, $D_{\alpha}^{*}(\rho_{2}\|\sigma_{2})=D_{2-\frac{1}{\alpha}}(\rho_{2}\|\sigma_{2})$. Hence, in the following, we only need to establish Eq. (14).
IV Proof of the achievability part
In this section, we focus on the proof of the achievability part of Eq. (14). The process is divided into three steps. In the first step, we construct an intermediate quantity $F(\rho_{1},\sigma_{1},\rho_{2},\sigma_{2},r)$ and derive its variational expression. In the second step, we prove that $F(\rho_{1},\sigma_{1},\rho_{2},\sigma_{2},r)$ is an upper bound of $E^{P}_{sc}(\rho_{1},\sigma_{1},\rho_{2},\sigma_{2},r)$. In the last step, we improve this upper bound to the correct form.
Definition 7
Let $\rho_{1},\sigma_{1}\in\mathcal{S}(\mathcal{H}_{1})$, $\rho_{2},\sigma_{2}\in\mathcal{S}(\mathcal{H}_{2})$ and $r\in\mathbb{R}$. The quantity $F(\rho_{1},\sigma_{1},\rho_{2},\sigma_{2},r)$ is defined as
$F(\rho_{1},\sigma_{1},\rho_{2},\sigma_{2},r):=\sup_{\frac{1}{2}<\alpha<1}\frac{1-\alpha}{\alpha}\big{\{}rD_{\alpha}^{\flat}(\rho_{2}\|\sigma_{2})-D_{\beta}^{\flat}(\rho_{1}\|\sigma_{1})\big{\}},$ (19)
where $\frac{1}{\alpha}+\frac{1}{\beta}=2$.
In the next proposition, we establish a variational expression for $F(\rho_{1},\sigma_{1},\rho_{2},\sigma_{2},r)$.
Proposition 8
For $\rho_{1},\sigma_{1}\in\mathcal{S}(\mathcal{H}_{1})$ that satisfy ${\operatorname{supp}}(\rho_{1})\subset{\operatorname{supp}}(\sigma_{1})$, $\rho_{2},\sigma_{2}\in\mathcal{S}(\mathcal{H}_{2})$ and $r\geq 0$, we have
$F(\rho_{1},\sigma_{1},\rho_{2},\sigma_{2},r)=\inf_{(\tau_{1},\tau_{2})\in\mathcal{S}_{\rho_{1}}\times\mathcal{S}_{\rho_{2}}}\{rD(\tau_{2}\|\rho_{2})+D(\tau_{1}\|\rho_{1})+|rD(\tau_{2}\|\sigma_{2})-D(\tau_{1}\|\sigma_{1})|^{+}\}.$
Proof. By Proposition 2(iii), we have
$\begin{split}&\sup_{\frac{1}{2}<\alpha<1}\frac{1-\alpha}{\alpha}\big{\{}rD_{\alpha}^{\flat}(\rho_{2}\|\sigma_{2})-D_{\beta}^{\flat}(\rho_{1}\|\sigma_{1})\big{\}}\\ =&\sup_{\frac{1}{2}<\alpha<1}\frac{1-\alpha}{\alpha}\big{\{}r\inf_{\tau_{2}\in\mathcal{S}_{\rho_{2}}}\big{(}D(\tau_{2}\|\sigma_{2})-\frac{\alpha}{\alpha-1}D(\tau_{2}\|\rho_{2})\big{)}-\sup_{\tau_{1}\in\mathcal{S}_{\rho_{1}}}\big{(}D(\tau_{1}\|\sigma_{1})-\frac{\beta}{\beta-1}D(\tau_{1}\|\rho_{1})\big{)}\big{\}}\\ =&\sup_{\frac{1}{2}<\alpha<1}\inf_{(\tau_{1},\tau_{2})\in\mathcal{S}_{\rho_{1}}\times\mathcal{S}_{\rho_{2}}}\frac{1-\alpha}{\alpha}\big{\{}rD(\tau_{2}\|\sigma_{2})-\frac{\alpha}{\alpha-1}rD(\tau_{2}\|\rho_{2})-D(\tau_{1}\|\sigma_{1})+\frac{\beta}{\beta-1}D(\tau_{1}\|\rho_{1})\big{\}}\\ =&\sup_{\frac{1}{2}<\alpha<1}\inf_{(\tau_{1},\tau_{2})\in\mathcal{S}_{\rho_{1}}\times\mathcal{S}_{\rho_{2}}}\big{\{}rD(\tau_{2}\|\rho_{2})+D(\tau_{1}\|\rho_{1})+\frac{1-\alpha}{\alpha}\big{(}rD(\tau_{2}\|\sigma_{2})-D(\tau_{1}\|\sigma_{1})\big{)}\big{\}}\\ =&\sup_{0<\delta<1}\inf_{(\tau_{1},\tau_{2})\in\mathcal{S}_{\rho_{1}}\times\mathcal{S}_{\rho_{2}}}\big{\{}rD(\tau_{2}\|\rho_{2})+D(\tau_{1}\|\rho_{1})+\delta\big{(}rD(\tau_{2}\|\sigma_{2})-D(\tau_{1}\|\sigma_{1})\big{)}\big{\}}\\ =&\inf_{(\tau_{1},\tau_{2})\in\mathcal{S}_{\rho_{1}}\times\mathcal{S}_{\rho_{2}}}\sup_{0<\delta<1}\big{\{}rD(\tau_{2}\|\rho_{2})+D(\tau_{1}\|\rho_{1})+\delta\big{(}rD(\tau_{2}\|\sigma_{2})-D(\tau_{1}\|\sigma_{1})\big{)}\big{\}}\\ =&\inf_{(\tau_{1},\tau_{2})\in\mathcal{S}_{\rho_{1}}\times\mathcal{S}_{\rho_{2}}}\{rD(\tau_{2}\|\rho_{2})+D(\tau_{1}\|\rho_{1})+|rD(\tau_{2}\|\sigma_{2})-D(\tau_{1}\|\sigma_{1})|^{+}\},\end{split}$ (20)
where the sixth line follows from Sion’s minimax theorem. The convexity of the function $(\tau_{1},\tau_{2})\mapsto rD(\tau_{2}\|\rho_{2})+D(\tau_{1}\|\rho_{1})+\delta\big{(}rD(\tau_{2}\|\sigma_{2})-D(\tau_{1}\|\sigma_{1})$ that the Sion’s minimax theorem requires is not obvious. To see this, we write
$\begin{split}&rD(\tau_{2}\|\rho_{2})+D(\tau_{1}\|\rho_{1})+\delta\big{(}rD(\tau_{2}\|\sigma_{2})-D(\tau_{1}\|\sigma_{1})\big{)}\\ =&rD(\tau_{2}\|\rho_{2})+\delta rD(\tau_{2}\|\sigma_{2})-(1-\delta)H(\tau_{1})+\operatorname{Tr}\tau_{1}(\delta\log\sigma_{1}-\log\rho_{1}).\end{split}$ (21)
[…]
Next, we show that $F(\rho_{1},\sigma_{1},\rho_{2},\sigma_{2},r)$ is an upper bound for $E^{P}_{sc}(\rho_{1},\sigma_{1},\rho_{2},\sigma_{2},r)$.
Lemma 9
For $\rho_{1},\sigma_{1}\in\mathcal{S}(\mathcal{H}_{1})$ satisfy ${\operatorname{supp}}(\rho_{1})\subset{\operatorname{supp}}(\sigma_{1})$, $\rho_{2},\sigma_{2}\in\mathcal{S}(\mathcal{H}_{2})$ and $r\geq 0$, we have
$E^{P}_{sc}(\rho_{1},\sigma_{1},\rho_{2},\sigma_{2},r)\leq F(\rho_{1},\sigma_{1},\rho_{2},\sigma_{2},r).$
In order to prove Lemma 9, we introduce
$\begin{split}&F_{1}(\rho_{1},\sigma_{1},\rho_{2},\sigma_{2},r):=\inf_{(\tau_{1},\tau_{2})\in\mathcal{F}_{1}}\{rD(\tau_{2}\|\rho_{2})+D(\tau_{1}\|\rho_{1})\}\\ &F_{2}(\rho_{1},\sigma_{1},\rho_{2},\sigma_{2},r):=\inf_{(\tau_{1},\tau_{2})\in\mathcal{F}_{2}}\{rD(\tau_{2}\|\rho_{2})+D(\tau_{1}\|\rho_{1})+rD(\tau_{2}\|\sigma_{2})-D(\tau_{1}\|\sigma_{1})\},\end{split}$ (22)
with
$\begin{split}&\mathcal{F}_{1}:=\big{\{}(\tau_{1},\tau_{2})~{}|~{}(\tau_{1},\tau_{2})\in\mathcal{S}_{\rho_{1}}\times\mathcal{S}_{\rho_{2}},r<\frac{D(\tau_{1}\|\sigma_{1})}{D(\tau_{2}\|\sigma_{2})}\big{\}},\\ &\mathcal{F}_{2}:=\big{\{}(\tau_{1},\tau_{2})~{}|~{}(\tau_{1},\tau_{2})\in\mathcal{S}_{\rho_{1}}\times\mathcal{S}_{\rho_{2}},r\geq\frac{D(\tau_{1}\|\sigma_{1})}{D(\tau_{2}\|\sigma_{2})}\big{\}}.\end{split}$
It is obviously seen from Proposition 8 that
$F(\rho_{1},\sigma_{1},\rho_{2},\sigma_{2},r)=\min\{F_{1}(\rho_{1},\sigma_{1},\rho_{2},\sigma_{2},r),F_{2}(\rho_{1},\sigma_{1},\rho_{2},\sigma_{2},r)\}.$ (23)
So it is suffices to prove that $F_{1}(\rho_{1},\sigma_{1},\rho_{2},\sigma_{2},r)$ and $F_{2}(\rho_{1},\sigma_{1},\rho_{2},\sigma_{2},r)$ are both the upper bounds for $E^{P}_{sc}(\rho_{1},\sigma_{1},\rho_{2},\sigma_{2},r)$. We accomplish the proof in the following Lemma 10 and Lemma 11 , respectively.
Lemma 10
For $\rho_{1},\sigma_{1}\in\mathcal{S}(\mathcal{H}_{1})$ satisfy ${\operatorname{supp}}(\rho_{1})\subset{\operatorname{supp}}(\sigma_{1})$, $\rho_{2},\sigma_{2}\in\mathcal{S}(\mathcal{H}_{2})$ and $r\geq 0$, we have
$E^{P}_{sc}(\rho_{1},\sigma_{1},\rho_{2},\sigma_{2},r)\leq F_{1}(\rho_{1},\sigma_{1},\rho_{2},\sigma_{2},r).$
Proof. According to the definition of $F_{1}(\rho_{1},\sigma_{1},\rho_{2},\sigma_{2},r)$, for any $\delta>0$, there exists $(\tau_{1},\tau_{2})\in\mathcal{S}_{\rho_{1}}\times\mathcal{S}_{\rho_{2}}$ such that
$r<\frac{D(\tau_{1}\|\sigma_{1})}{D(\tau_{2}\|\sigma_{2})}$ (24)
and
$rD(\tau_{2}\|\rho_{2})+D(\tau_{1}\|\rho_{1})\leq F_{1}(\rho_{1},\sigma_{1},\rho_{2},\sigma_{2},r)+\delta.$ (25)
Eq. (24) implies that $r$ is an asymptotically achievable rate for the transformation from $(\tau_{1},\sigma_{1})$ to $(\tau_{2},\sigma_{2})$. Hence, there exists a sequence of quantum channels $\{\mathcal{E}_{n}\}_{n\in\mathbb{N}}$ such that
$\begin{split}&\mathcal{E}_{n}(\sigma_{1}^{\otimes n})=\sigma_{2}^{\otimes nr},\\ &\lim_{n\rightarrow\infty}\|\mathcal{E}_{n}(\tau_{1}^{\otimes n})-\tau_{2}^{\otimes nr}\|_{1}=0.\end{split}$ (26)
Now, we let $t=4(rD(\tau_{2}\|\rho_{2})+D(\tau_{1}\|\rho_{1}))$, $\epsilon_{n}=2^{-nt}$, $C=\max\{1-r\log\lambda_{\rm{min}}(\rho_{2}),t+r\log|\mathcal{H}_{2}|\}$, where $\lambda_{\rm{min}}(\rho_{2})$ is the smallest non-zero eigenvalue of $\rho_{2}$ and $\rho_{2}^{n}=(1-\epsilon_{n})\rho_{2}^{\otimes nr}+\epsilon_{n}\pi_{n}$, where $\pi_{n}$ is the maximally mixed state onto the subspace ${\operatorname{supp}}(\rho_{2}^{\otimes nr})^{\perp}$. For sufficiently large $n$, $\lambda_{\rm{min}}(\rho_{2}^{n})\geq 2^{-nC}$ and for such $n$, $F(\rho_{2}^{n},\mathcal{E}_{n}(\rho_{1}^{\otimes n}))$ can be evaluated as follows.
$\begin{split}&-\log F^{2}(\rho_{2}^{n},\mathcal{E}_{n}(\rho_{1}^{\otimes n}))\\ \leq&D(\mathcal{E}_{n}(\tau_{1}^{\otimes n})\|\mathcal{E}_{n}(\rho_{1}^{\otimes n}))+D(\mathcal{E}_{n}(\tau_{1}^{\otimes n})\|\rho_{2}^{n})\\ \leq&nD(\tau_{1}\|\rho_{1})+D(\mathcal{E}_{n}(\tau_{1}^{\otimes n})\|\rho_{2}^{n})\\ =&nD(\tau_{1}\|\rho_{1})-H(\mathcal{E}_{n}(\tau_{1}^{\otimes n}))-\operatorname{Tr}\mathcal{E}_{n}(\tau_{1}^{\otimes n})\log\rho_{2}^{n}+H(\tau_{2}^{\otimes nr})+\operatorname{Tr}\tau_{2}^{\otimes nr}\log\rho_{2}^{n}+D(\tau_{2}^{\otimes nr}\|\rho_{2}^{n})\\ =&nD(\tau_{1}\|\rho_{1})-H(\mathcal{E}_{n}(\tau_{1}^{\otimes n}))+H(\tau_{2}^{\otimes nr})+\operatorname{Tr}\big{(}\tau_{2}^{\otimes nr}-\mathcal{E}_{n}(\tau_{1}^{\otimes n})\big{)}\log\rho_{2}^{n}+D(\tau_{2}^{\otimes nr}\|\rho_{2}^{n})\\ \leq&nD(\tau_{1}\|\rho_{1})+\frac{1}{2}\|\mathcal{E}_{n}(\tau_{1}^{\otimes n})-\tau_{2}^{\otimes nr}\|_{1}\log|\mathcal{H}_{2}|^{nr}+h(\frac{1}{2}\|\mathcal{E}_{n}(\tau_{1}^{\otimes n})-\tau_{2}^{\otimes nr}\|_{1})\\ &+\|\mathcal{E}_{n}(\tau_{1}^{\otimes n})-\tau_{2}^{\otimes nr}\|_{1}\|\log\rho_{2}^{n}\|_{\infty}+D(\tau_{2}^{\otimes nr}\|\rho_{2}^{n})\\ \leq&nD(\tau_{1}\|\rho_{1})+\frac{1}{2}\|\mathcal{E}_{n}(\tau_{1}^{\otimes n})-\tau_{2}^{\otimes nr}\|_{1}\log|\mathcal{H}_{2}|^{nr}+h(\frac{1}{2}\|\mathcal{E}_{n}(\tau_{1}^{\otimes n})-\tau_{2}^{\otimes nr}\|_{1})\\ &+nC\|\mathcal{E}_{n}(\tau_{1}^{\otimes n})-\tau_{2}^{\otimes nr}\|_{1}+D(\tau_{2}^{\otimes nr}\|\rho_{2}^{n})\\ \leq&nD(\tau_{1}\|\rho_{1})+\frac{1}{2}\|\mathcal{E}_{n}(\tau_{1}^{\otimes n})-\tau_{2}^{\otimes nr}\|_{1}\log|\mathcal{H}_{2}|^{nr}+h(\frac{1}{2}\|\mathcal{E}_{n}(\tau_{1}^{\otimes n})-\tau_{2}^{\otimes nr}\|_{1})\\ &+nC\|\mathcal{E}_{n}(\tau_{1}^{\otimes n})-\tau_{2}^{\otimes nr}\|_{1}+nrD(\tau_{2}\|\rho_{2})-\log(1-\epsilon_{n}),\end{split}$ (27)
where the second line comes from Lemma 15 in Appendix, the third line is due to the data processing inequality, the sixth line follows from the Fannes-Audenaert inequality and the operator Hölder inequality. Dividing $n$ in both sides of Eq. (27) and taking the limit of $n$, we have
$\lim_{n\rightarrow\infty}-\frac{\log F^{2}(\rho_{2}^{n},\mathcal{E}_{n}(\rho_{1}^{\otimes n}))}{n}\leq D(\tau_{1}\|\rho_{1})+rD(\tau_{2}\|\rho_{2}).$ (28)
By Lemma 13 in Appendix, the following inequality holds:
$\begin{split}F^{2}(\mathcal{E}_{n}(\rho_{1}^{\otimes n}),\rho_{2}^{\otimes nr})&\geq F^{2}(\rho_{2}^{n},\mathcal{E}_{n}(\rho_{1}^{\otimes n}))-\sqrt{\|\rho_{2}^{n}-\rho_{2}^{\otimes nr}\|_{1}}\\ &=F^{2}(\rho_{2}^{n},\mathcal{E}_{n}(\rho_{1}^{\otimes n}))-\sqrt{2\epsilon_{n}}.\end{split}$ (29)
Eq. (28), Eq. (29) and the definition of $\epsilon_{n}$ gives that
$\lim_{n\rightarrow\infty}-\frac{\log F^{2}(\rho_{2}^{\otimes nr},\mathcal{E}_{n}(\rho_{1}^{\otimes n}))}{n}\leq D(\tau_{1}\|\rho_{1})+rD(\tau_{2}\|\rho_{2})\leq F_{1}(\rho_{1},\sigma_{1},\rho_{2},\sigma_{2},r)+\delta.$ (30)
Eq. (30) and Remark 4 implies that
$E^{P}_{sc}(\rho_{1},\sigma_{1},\rho_{2},\sigma_{2},r)\leq F_{1}(\rho_{1},\sigma_{1},\rho_{2},\sigma_{2},r)+\delta.$ (31)
[…]
Lemma 11
For $\rho_{1},\sigma_{1}\in\mathcal{S}(\mathcal{H}_{1})$ satisfy ${\operatorname{supp}}(\rho_{1})\subset{\operatorname{supp}}(\sigma_{1})$, $\rho_{2},\sigma_{2}\in\mathcal{S}(\mathcal{H}_{2})$ and $r\geq 0$, we have
$E^{P}_{sc}(\rho_{1},\sigma_{1},\rho_{2},\sigma_{2},r)\leq F_{2}(\rho_{1},\sigma_{1},\rho_{2},\sigma_{2},r).$
Proof. According to the definition of $F_{2}(\rho_{1},\sigma_{1},\rho_{2},\sigma_{2},r)$, there exists $(\tau_{1},\tau_{2})\in\mathcal{S}_{\rho_{1}}\times\mathcal{S}_{\rho_{2}}$ such that
$r\geq\frac{D(\tau_{1}\|\sigma_{1})}{D(\tau_{2}\|\sigma_{2})}$ (32)
and
$F_{2}(\rho_{1},\sigma_{1},\rho_{2},\sigma_{2},r)=rD(\tau_{2}\|\rho_{2})+D(\tau_{1}\|\rho_{1})+rD(\tau_{2}\|\sigma_{2})-D(\tau_{1}\|\sigma_{1}).$ (33)
We let $r_{1}=\frac{D(\tau_{1}\|\sigma_{1})}{D(\tau_{2}\|\sigma_{2})}-\delta$, where $\delta>0$ is a constant. From the proof of Lemma 10, we know that there exists a sequence of quantum channels $\{\mathcal{E}_{n}\}_{n\in\mathbb{N}}$ such that
$\mathcal{E}_{n}(\sigma_{1}^{\otimes n})=\sigma_{2}^{\otimes nr_{1}}$ (34)
and
$\lim_{n\rightarrow\infty}-\frac{\log F^{2}(\rho_{2}^{\otimes nr_{1}},\mathcal{E}_{n}(\rho_{1}^{\otimes n}))}{n}\leq D(\tau_{1}\|\rho_{1})+r_{1}D(\tau_{2}\|\rho_{2}).$ (35)
Now, we construct a sequence of quantum channels $\{\Phi_{n}\}_{n\in\mathbb{N}}$ as
$\Phi_{n}(X)=\mathcal{E}_{n}(X)\otimes\sigma_{2}^{\otimes n(r-r_{1})}$
It is easy to verify that
$\Phi_{n}(\sigma_{1}^{\otimes n})=\sigma_{2}^{\otimes nr}$ (36)
and
$\begin{split}&\lim_{n\rightarrow\infty}-\frac{\log F^{2}(\rho_{2}^{\otimes nr},\Phi_{n}(\rho_{1}^{\otimes n}))}{n}\\ =&\lim_{n\rightarrow\infty}-\big{\{}\frac{\log F^{2}(\rho_{2}^{\otimes nr_{1}},\mathcal{E}_{n}(\rho_{1}^{\otimes n}))}{n}+\frac{\log F^{2}(\rho_{2}^{\otimes n(r-r_{1})},\sigma_{2}^{\otimes n(r-r_{1})})}{n}\big{\}}\\ \leq&D(\tau_{1}\|\rho_{1})+r_{1}D(\tau_{2}\|\rho_{2})-(r-r_{1})\log F^{2}(\rho_{2},\sigma_{2})\\ \leq&D(\tau_{1}\|\rho_{1})+r_{1}D(\tau_{2}\|\rho_{2})+(r-r_{1})\{D(\tau_{2}\|\rho_{2})+D(\tau_{2}\|\sigma_{2})\}\\ =&D(\tau_{1}\|\rho_{1})+rD(\tau_{2}\|\rho_{2})+rD(\tau_{2}\|\sigma_{2})-D(\tau_{1}\|\sigma_{1})+\delta D(\tau_{2}\|\sigma_{2}),\end{split}$ (37)
where in second inequality, we use Lemma 15 in Appendix. Eq. (33), Eq. (36) and Eq. (37) imply that
$E^{P}_{sc}(\rho_{1},\sigma_{1},\rho_{2},\sigma_{2},r)\leq F_{2}(\rho_{1},\sigma_{1},\rho_{2},\sigma_{2},r)+\delta D(\tau_{2}\|\sigma_{2}).$ (38)
[…]
In the last step, we improve the the upper bound in Lemma 9 to the correct form to finish the proof of the achievability part of Eq. (14).
Proof of the achievability part of Eq. (14). If ${\operatorname{supp}}(\rho_{1})\subseteq{\operatorname{supp}}(\sigma_{1})$ does not hold, it is obvious that all terms in Eq. (14) are equal to 00 and the assertion holds trivially. Hence, for the rest we assume that ${\operatorname{supp}}(\rho_{1})\subseteq{\operatorname{supp}}(\sigma_{1})$. We first fix an integer $m$ and let $\rho_{1}^{m}=\mathcal{E}_{\sigma_{1}^{\otimes m}}(\rho_{1}^{\otimes m})$, $\rho_{2}^{m}=\mathcal{E}_{\sigma_{2}^{\otimes m}}(\rho_{2}^{\otimes m})$. From Lemma 9 , we know that there exists a sequence of quantum channels $\{\Phi_{k}\}_{k\in\mathbb{N}}$ such that
$\Phi_{k}(\sigma_{1}^{\otimes mk})=\sigma_{2}^{\otimes mkr}$ (39)
and
$\lim_{k\rightarrow\infty}-\frac{\log F^{2}({\rho^{m}_{2}}^{\otimes kr},\Phi_{k}({\rho^{m}_{1}}^{\otimes k}))}{k}\leq\sup_{\frac{1}{2}<\alpha<1}\frac{1-\alpha}{\alpha}\big{\{}rD_{\alpha}^{\flat}(\rho^{m}_{2}\|\sigma^{\otimes m}_{2})-D_{\beta}^{\flat}(\rho^{m}_{1}\|\sigma^{\otimes m}_{1})\big{\}}.$ (40)
Lemma 14 in Appendix and Eq. (40) imply that
$\begin{split}&\lim_{k\rightarrow\infty}-\frac{\log F^{2}({\rho_{2}}^{\otimes mkr},\mathcal{E}_{\sigma_{2}^{\otimes m}}^{\otimes kr}\Phi_{k}({\rho^{m}_{1}}^{\otimes k}))}{k}\\ \leq&\sup_{\frac{1}{2}<\alpha<1}\frac{1-\alpha}{\alpha}\big{\{}rD_{\alpha}^{\flat}(\rho^{m}_{2}\|\sigma^{\otimes m}_{2})-D_{\beta}^{\flat}(\rho^{m}_{1}\|\sigma^{\otimes m}_{1})\big{\}}+r\log v(\sigma_{2}^{\otimes m}).\end{split}$ (41)
Now, for any integer $n$, we can write $n=km+l$, where $10\leq l\leq m-$ 1 and we construct a quantum channel $\mathcal{E}_{n}$ as
$\mathcal{E}_{n}(X)=\mathcal{E}_{\sigma_{2}^{\otimes m}}^{\otimes kr}\Phi_{k}\mathcal{E}_{\sigma_{1}^{\otimes m}}^{\otimes k}\operatorname{Tr}_{l}(X)\otimes\sigma_{2}^{\otimes lr},$
where $\operatorname{Tr}_{l}$ denotes taking the partial trace over the last $l$ systems.
From Eq. (39) and Eq. (41), it is easy to see that
$\mathcal{E}_{n}(\sigma_{1}^{\otimes n})=\sigma_{2}^{\otimes nr}$ (42)
and
$\begin{split}&\lim_{n\rightarrow\infty}-\frac{\log F^{2}({\rho_{2}}^{\otimes nr},\mathcal{E}_{n}(\rho_{1}^{\otimes n}))}{n}\\ =&\lim_{k\rightarrow\infty}-\frac{\log F^{2}({\rho_{2}}^{\otimes mkr},\mathcal{E}_{\sigma_{2}^{\otimes m}}^{\otimes kr}\Phi_{k}({\rho^{m}_{1}}^{\otimes k}))F^{2}(\rho_{2}^{\otimes lr},\sigma_{2}^{\otimes lr})}{mk+l}\\ \leq&\sup_{\frac{1}{2}<\alpha<1}\frac{1-\alpha}{\alpha}\big{\{}r\frac{D_{\alpha}^{\flat}(\rho^{m}_{2}\|\sigma^{\otimes m}_{2})}{m}-\frac{D_{\beta}^{\flat}(\rho^{m}_{1}\|\sigma^{\otimes m}_{1})}{m}\big{\}}+r\frac{\log v(\sigma_{2}^{\otimes m})}{m}.\end{split}$ (43)
Eq. (42) and Eq. (43) give that
$\begin{split}&E^{P}_{sc}(\rho_{1},\sigma_{1},\rho_{2},\sigma_{2},r)\\ \leq&\sup_{\frac{1}{2}<\alpha<1}\frac{1-\alpha}{\alpha}\big{\{}r\frac{D_{\alpha}^{\flat}(\rho^{m}_{2}\|\sigma^{\otimes m}_{2})}{m}-\frac{D_{\beta}^{\flat}(\rho^{m}_{1}\|\sigma^{\otimes m}_{1})}{m}\big{\}}+r\frac{\log v(\sigma_{2}^{\otimes m})}{m}\\ =&\sup_{\frac{1}{2}<\alpha<1}\frac{1-\alpha}{\alpha}\big{\{}r\frac{D_{\alpha}^{*}(\rho^{m}_{2}\|\sigma^{\otimes m}_{2})}{m}-\frac{D_{\beta}^{*}(\rho^{m}_{1}\|\sigma^{\otimes m}_{1})}{m}\big{\}}+r\frac{\log v(\sigma_{2}^{\otimes m})}{m}\\ \leq&\sup_{\frac{1}{2}<\alpha<1}\frac{1-\alpha}{\alpha}\big{\{}rD_{\alpha}^{*}(\rho_{2}\|\sigma_{2})-D_{\beta}^{*}(\rho_{1}\|\sigma_{1})+2\frac{v(\sigma_{1}^{\otimes m})}{m}\big{\}}+r\frac{\log v(\sigma_{2}^{\otimes m})}{m}\\ \leq&\sup_{\frac{1}{2}<\alpha<1}\frac{1-\alpha}{\alpha}\big{\{}rD_{\alpha}^{*}(\rho_{2}\|\sigma_{2})-D_{\beta}^{*}(\rho_{1}\|\sigma_{1})\big{\}}+2\frac{v(\sigma_{1}^{\otimes m})}{m}+r\frac{\log v(\sigma_{2}^{\otimes m})}{m}\\ =&\sup_{\frac{1}{2}\leq\alpha\leq 1}\frac{1-\alpha}{\alpha}\big{\{}rD_{\alpha}^{*}(\rho_{2}\|\sigma_{2})-D_{\frac{\alpha}{2\alpha-1}}^{*}(\rho_{1}\|\sigma_{1})\big{\}}+2\frac{v(\sigma_{1}^{\otimes m})}{m}+r\frac{\log v(\sigma_{2}^{\otimes m})}{m},\end{split}$ (44)
[…]
V Proof of the optimality part
In this section, we prove the optimality part of Eq. (14). The proof is essentially from a easy application of the operator Hölder inequality.
Proof of the optimality part of Eq. (14). If ${\operatorname{supp}}(\rho_{1})\subseteq{\operatorname{supp}}(\sigma_{1})$ does not hold, it is obvious that all terms in Eq. (14) are equal to 00 and the assertion holds trivially. Hence, for the rest we assume that ${\operatorname{supp}}(\rho_{1})\subseteq{\operatorname{supp}}(\sigma_{1})$. By Lemma 16 in Appendix, for any quantum channel $\mathcal{E}_{n}$ such that $\mathcal{E}_{n}(\sigma_{1}^{\otimes n})=\sigma_{2}^{\otimes nr}$ and any $\alpha\in(\frac{1}{2},1)$, $\beta=\frac{\alpha}{2\alpha-1}$, we have
$\begin{split}&\frac{2\alpha}{1-\alpha}\log F\big{(}\mathcal{E}_{n}(\rho_{1}^{\otimes n}),\rho_{2}^{\otimes nr}\big{)}\\ \leq&-D_{\alpha}^{*}(\rho_{2}^{\otimes nr}\|\sigma_{2}^{\otimes nr})+D^{*}_{\beta}(\mathcal{E}_{n}(\rho_{1}^{\otimes n})\|\sigma_{2}^{\otimes nr})\\ =&-D_{\alpha}^{*}(\rho_{2}^{\otimes nr}\|\sigma_{2}^{\otimes nr})+D^{*}_{\beta}(\mathcal{E}_{n}(\rho_{1}^{\otimes n})\|\mathcal{E}_{n}(\sigma_{1}^{\otimes n}))\\ \leq&-nrD_{\alpha}^{*}(\rho_{2}\|\sigma_{2})+nD_{\beta}^{*}(\rho_{1}\|\sigma_{1}).\end{split}$ (45)
Eq. (45) implies that
$\begin{split}E^{P}_{sc}(\rho_{1},\sigma_{1},\rho_{2},\sigma_{2},r)&=\lim_{n\rightarrow\infty}\frac{-1}{n}\log\max_{\mathcal{E}_{n}:\mathcal{E}_{n}(\sigma_{1}^{\otimes n})=\sigma_{2}^{\otimes nr}}F^{2}\big{(}\mathcal{E}_{n}(\rho_{1}^{\otimes n}),\rho_{2}^{\otimes nr}\big{)}\\ &\geq\frac{1-\alpha}{\alpha}\big{\{}rD_{\alpha}^{*}(\rho_{2}\|\sigma_{2})-D_{\beta}^{*}(\rho_{1}\|\sigma_{1})\big{\}}.\end{split}$ (46)
Noticing that Eq. (46) holds for any $\alpha\in(\frac{1}{2},1)$, we have
$\begin{split}E^{P}_{sc}(\rho_{1},\sigma_{1},\rho_{2},\sigma_{2},r)&\geq\sup_{\frac{1}{2}<\alpha<1}\frac{1-\alpha}{\alpha}\big{\{}rD_{\alpha}^{*}(\rho_{2}\|\sigma_{2})-D_{\beta}^{*}(\rho_{1}\|\sigma_{1})\big{\}}\\ &=\sup_{\frac{1}{2}\leq\alpha\leq 1}\frac{1-\alpha}{\alpha}\big{\{}rD_{\alpha}^{*}(\rho_{2}\|\sigma_{2})-D_{\beta}^{*}(\rho_{1}\|\sigma_{1})\big{\}}\\ &=\sup_{\frac{1}{2}\leq\alpha\leq 1}\frac{1-\alpha}{\alpha}\big{\{}rD_{\alpha}^{*}(\rho_{2}\|\sigma_{2})-D_{\frac{\alpha}{2\alpha-1}}^{*}(\rho_{1}\|\sigma_{1})\big{\}}.\end{split}$ (47)
[…]
VI Discussion
In this work, we determine the exact strong converse exponent of quantum dichotomies based on the purified distance. It is characterized by the sandwiched Rényi divergence $D_{\beta}(\rho_{1}\|\sigma_{1})$ of the initial quantum dichotomy with $\beta\in(1,\infty)$ and the sandwiched Rényi divergence $D_{\alpha}(\rho_{2}\|\sigma_{2})$ of the target quantum dichotomy $(\rho_{2},\sigma_{2})$ with $\alpha\in(\frac{1}{2},1)$. Hence, we find a new type of multi-variate quantum Rényi information measure involving four quantum states — with an exact operational meaning. The combination of this result and an improved Fuchs–van de Graaf inequality also directly leads to the strong converse exponent based on the trace distance when $\rho_{2}$ is a pure state. We further believe that this formula still holds when $\rho_{2}$ is a general quantum state, but leave it as an open question. While recent results have established lower bounds for the strong converse exponents under the trace distance for some tasks Wilde2022distinguishability ; SalzmannDatta2022total ; SGC2022optimal , deriving a matching upper bound seems a difficult problem. The improved Fuchs–van de Graaf inequality provides a new idea to solve this kind of problem when dealing with pure states. In this case, the strong converse exponent under the trace distance can be transformed into that under the purified distance. For determining the strong converse exponent under the purified distance, there are then already a few techniques developed MosonyiOgawa2017strong ; LiYao2024operational .
Acknowledgments
MB and YY acknowledge funding by the European Research Council (ERC Grant Agreement No. 948139) and MB acknowledges support from the Excellence Cluster - Matter and Light for Quantum Computing (ML4Q).
VII Appendix
The following lemma is an improved Fuchs–van de Graaf inequality.
Lemma 12
For any pure state $\varphi$ and subnormalized state $\rho$, we have
$P(\rho,\varphi)\leq\sqrt{d(\varphi,\rho)}.$ (48)
Proof. By direct calculation, we have
$d(\varphi,\rho)=\operatorname{Tr}(\varphi-\rho)_{+}\geq\operatorname{Tr}(\varphi-\rho)\varphi=1-\operatorname{Tr}\rho\varphi=1-F^{2}(\rho,\varphi)=P^{2}(\rho,\varphi).$ (49)
[…]
Lemma 13
For $\rho,\sigma,\tau\in\mathcal{S}(\mathcal{H})$, we have
$F^{2}(\tau,\sigma)\geq F^{2}(\rho,\sigma)-\sqrt{\|\rho-\tau\|_{1}}.$ (50)
Proof. By Ulhmann’s theorem Uhlmann1976transition , there exists purified states $\phi_{\rho}$, $\phi_{\sigma}$ and $\phi_{\tau}$ of $\rho$, $\sigma$ and $\tau$ such that
$F(\rho,\sigma)=F(\phi_{\rho},\phi_{\sigma}),F(\rho,\tau)=F(\phi_{\rho},\phi_{\tau}).$
By direct calculation, we have
$\begin{split}F^{2}(\tau,\sigma)&\geq F^{2}(\phi_{\tau},\phi_{\sigma})\\ &=\operatorname{Tr}\phi_{\tau}\phi_{\sigma}\\ &=\operatorname{Tr}\phi_{\sigma}\phi_{\rho}+\operatorname{Tr}\phi_{\sigma}(\phi_{\tau}-\phi_{\rho})\\ &\geq\operatorname{Tr}\phi_{\sigma}\phi_{\rho}-\operatorname{Tr}(\phi_{\rho}-\phi_{\tau})_{+}\\ &=F^{2}(\rho,\sigma)-P(\phi_{\rho},\phi_{\tau})\\ &=F^{2}(\rho,\sigma)-P(\rho,\tau)\\ &\geq F^{2}(\rho,\sigma)-\sqrt{\|\rho-\tau\|_{1}}.\end{split}$ (51)
[…]
The following three lemmas are used in our proofs and they have been proved in LiYao2024operational .
Lemma 14
Let $\rho,\sigma\in\mathcal{S}(\mathcal{H})$, and let $\mathcal{H}=\bigoplus_{i\in\mathcal{I}}\mathcal{H}_{i}$ decompose into a set of mutually orthogonal subspaces $\{\mathcal{H}_{i}\}_{i\in\mathcal{I}}$. Suppose that $\sigma=\sum\limits_{i\in\mathcal{I}}\sigma_{i}$ with ${\operatorname{supp}}(\sigma_{i})\subseteq\mathcal{H}_{i}$. Then
$F\left(\sum_{i\in\mathcal{I}}\Pi_{i}\rho\Pi_{i},\sigma\right)\leq\sqrt{|\mathcal{I}|}F(\rho,\sigma),$ (52)
where $\Pi_{i}$ is the projection onto $\mathcal{H}_{i}$.
Lemma 15
Let $\rho,\sigma,\tau\in\mathcal{S}(\mathcal{H})$ be any quantum states. Then we have
$-\log F^{2}(\rho,\sigma)\leq D(\tau\|\rho)+D(\tau\|\sigma).$ (53)
Lemma 16
Let $\rho,\sigma\in\mathcal{S}(\mathcal{H})$ and $\tau\in\mathcal{P}(\mathcal{H})$, and suppose ${\operatorname{supp}}(\sigma)\not\perp{\operatorname{supp}}(\tau)$. Fix $\alpha\in(\frac{1}{2},1)$ and $\beta\in(1,+\infty)$ such that $\frac{1}{\alpha}+\frac{1}{\beta}=2$. Then
$\frac{2\alpha}{1-\alpha}\log F(\rho,\sigma)\leq D^{*}_{\beta}(\rho\|\tau)-D^{*}_{\alpha}(\sigma\|\tau).$ (54)